OpenStudy (anonymous):

FAN AND MEDAL!!!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). A) y = -1/3x - 6 B) y = -1/3x - 14 C) y = 3x - 6 D) y = 3x - 14

OpenStudy (anonymous):

@DUDE..IM..A..DUCK

OpenStudy (anonymous):

quack

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

idk if im Graphing right because none of them went through (4, -2)

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

-_- im not gonna fan you

OpenStudy (anonymous):

luls

OpenStudy (anonymous):

why are you doing this? im trying to do school work here and your just trolling. find someone else or help me with the problem.

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

eq. of any line parallel to line 3 x-y=-3 is 3x-y=c it passes through (4,-2) 3*4-(-2)=c c=12+2=14 eq. of line is 3 x-y=14

OpenStudy (anonymous):

thank you, I understand now

OpenStudy (anonymous):

yw